OpenStudy (anonymous):

find the unit vector in the direction of (7,-24)

OpenStudy (zzr0ck3r):

divide that vector by the magnitude of that vector \((\frac{7}{\sqrt{7^2+24^2}{}},\frac{-24}{\sqrt{7^2+24^2}})\)